If you invest $1000 now at an interest rate of 5% compound ed annually, how many annual withdrawals of $200 could you make until the fund is exhausted (closest to zero)?
Answer #1

Annual withdrawals = $200. Amount invested = $1000. Rate = 5%, Time = n

This information suggests that

present value = annual withdrawal x (P/A, i%, n)

1000 = 200 x (P/A, 5%, n)

(P/A, 5%, n) = 5

For n = 6, (P/A, 5%, n) = 5.0757 and for n = 5, (P/A, 5%, n) = 4.3295.

Hence n should be 5 and we can make 5 annual withdrawals of $200 each with $1000 deposited now.